bus a takes 21 minutes to complete its route. bus b takes 35 minutes to complete its route. at what time will they arrive at the bus station at the same time?
Answer by stanbon(75887)

bus a takes 21 minutes to complete its route. bus b takes 35 minutes to complete its route. at what time will they arrive at the bus station at the same time?
-------
Find the least common multiple of 21 and 35:
Prime Factor form:
21 = 3*7
35 = 5*7
-----
least common multiple = 3*5*7 = 105
----
Answer:: 105 minutes
===========
Cheers,
Stan H.